# Towards the quark mass dependence of $T_{cc}^{+}$ from lattice QCD

### v1

@@ VII Conclusions
We employ two types of analysis. The first procedure utilises an effective range expansion for the energy dependence of the $DD^{*}$ scattering amplitude in the elastic region above threshold. Ignoring potentially nonnegligible left-hand cut effects and naively extrapolating the resultant energy dependence below the threshold renders a virtual state along the real axis below threshold for all the heavy-quark masses (corresponding to $DD^{*}$ reduced masses in the range $m_{r}\simeq 0.9-1.2$ GeV). The observed quark mass dependence of the ERE parameters and the resultant pole are summarised in Figs. 5-7. The virtual pole approaches the threshold as the heavy-quark mass is increased and is expected to become a bound state, related to $T_{cc}^{+}$, at $m_{r}^{\rm crit}=1.35(^{+17}_{-13})$ GeV.
@@
Our analysis favours weak overall dependence of the $D$-$D^{*}$ interaction on the heavy-quark mass and stronger bound systems for heavier constituents. This is in agreement with general expectations that the binding energy of a doubly-heavy system mainly scales with the kinetic term that is inversely proportional to the heavy mass. The estimated value of the critical reduced mass from Eq. (LABEL:critERE), when the virtual state converts to a bound state, is very close to the reduced mass of the physical $B^{*}D$ system $m_{r}^{\rm ph}(B^{*}D)\simeq 1.38$ GeV relevant for the $T_{bc}$ state with $J^{P}=1^{+}$. This observation hints towards the existence of the $T_{bc}$ pole (bound or virtual) very near the $B^{*}D$ threshold. However, this insight needs to be taken with caution since our study explicitly considers two heavy quarks of the same flavour and relies on the effective range expansion, which is not justified in the presence of a near-by left-hand cut. A similar estimate based on the effective theory approach is ambiguous. On the one hand, the dependence of $c_{0}$ and $c_{2}$ on the spin average $D$ meson mass in Fig. 10 is consistent (in the sense of a $\chi^{2}$ fit) with $1/\bar{m}_{D}$ and $1/\bar{m}_{D}^{2}$, respectively. [^9] We note that, for the $X(3872)$, the $1/\bar{m}_{D}$ mass scaling of the leading-order contact term was proposed in Ref. [53] based on demanding a proper power counting for a heavy-heavy system. However, according to the claim of Ref. [54], such a dependence (as well as any other) cannot be derived directly from effective field theory. On the other hand, in both plots, the most right point corresponding to the largest heavy-quark mass has a large uncertainty and is basically ignored by the fit. Under these circumstances, the behaviour of the contact terms cannot be reliably extrapolated to larger values of $\bar{m}_{D}$ consistent with the $T_{bc}$ case. Thus, at the present stage we cannot make a definite quantitative conclusion concerning the nature of the $T_{bc}$. Direct lattice studies of this state have not reached complete consensus yet as to whether it would be a shallow bound state [25, 26] or not [27, 28]. This intriguing tetraquark is being experimentally searched for by LHCb.
@@ Acknowledgements.
The authors would like to thank M. Doering, M.-L. Du, A. Filin, C. Hanhart, M. Hansen, B. Kubis, L. Leskovec, B. Mevlja, M. Mikhasenko, and A. Milstein for valuable discussions. A.N. and S.P. are supported by the Slovenian Research Agency (research core Funding No. P1-0035). A.N. also acknowledges support from the CAS President’s International Fellowship Initiative (PIFI) (Grant No. 2024PVA0004). M.P. gratefully acknowledges support from the Department of Science and Technology, India, SERB Start-up Research Grant No. SRG/2023/001235 and Department of Atomic Energy, India. We thank our colleagues in CLS for the joint effort in the generation of the gauge field ensembles which form a basis for the computation. We use the multigrid solver of Refs. [55, 56, 57, 58] for the inversion of the Dirac operator. Our code implementing distillation is written within the framework of the Chroma software package [59]. The simulations were performed on the Regensburg Athene2 cluster. We thank the authors of Ref. [60] for making the TwoHadronsInBox package public.
@@ Appendix A Discussion of the $DD^{*}$ interaction potential
[$*{}^{*}$ 0.348 −$-3.23\pm 0.39$ $6.37\pm 1.48$−$\simeq\leavevmode\nobreak\ -15\pm 19\leavevmode\nobreak\ i$$-2.00\pm 0.47$ 0
@@ Appendix B Approach to solving Lippmann–Schwinger equation (21)
Then we find (with the subscripts ${}_{I}$ and ${}_{II}$ for the first and second Riemann sheet, respectively) that, across the unitary cut,

### v3

@@ VII Conclusions
We employ two types of analysis. The first procedure utilises an effective range expansion for the energy dependence of the $DD^{*}$ scattering amplitude in the elastic region above threshold. Ignoring potentially nonnegligible left-hand cut effects and naively extrapolating the resultant energy dependence below the threshold renders a virtual state along the real axis below threshold for all the heavy-quark masses (corresponding to $DD^{*}$ reduced masses in the range $m_{r}\simeq 0.9-1.2$ GeV). The observed quark mass dependence of the effective range parameters and the resultant pole are summarised in Figs. 5-7. The virtual pole approaches the threshold as the heavy-quark mass is increased and is expected to become a bound state, related to $T_{cc}^{+}$, at $m_{r}^{\rm crit}=1.35(^{+17}_{-13})$ GeV.
@@
Our analysis favours weak overall dependence of the $D$-$D^{*}$ interaction on the heavy-quark mass and stronger bound systems for heavier constituents. This is in agreement with general expectations that the binding energy of a doubly-heavy system mainly scales with the kinetic term that is inversely proportional to the heavy mass. The estimated value of the critical reduced mass from Eq. (LABEL:critERE), when the virtual state converts to a bound state, is very close to the reduced mass of the physical $B^{*}D$ system $m_{r}^{\rm ph}(B^{*}D)\simeq 1.38$ GeV relevant for the $T_{bc}$ state with $J^{P}=1^{+}$. This observation hints towards the existence of the $T_{bc}$ pole (bound or virtual) very near the $B^{*}D$ threshold. However, this insight needs to be taken with caution since our study explicitly considers two heavy quarks of the same flavour and relies on the effective range expansion, which is not justified in the presence of a near-by left-hand cut. A similar estimate based on the effective theory approach is ambiguous. On the one hand, the dependence of $c_{0}$ and $c_{2}$ on the spin average $D$ meson mass in Fig. 10 is consistent (in the sense of a $\chi^{2}$ fit) with $1/\bar{m}_{D}$ and $1/\bar{m}_{D}^{2}$, respectively. [^11] We note that, for the $X(3872)$, the $1/\bar{m}_{D}$ mass scaling of the leading-order contact term was proposed in Ref. [54] based on demanding a proper power counting for a heavy-heavy system. However, according to the claim of Ref. [55], such a dependence (as well as any other) cannot be derived directly from effective field theory. On the other hand, in both plots, the most right point corresponding to the largest heavy-quark mass has a large uncertainty and is basically ignored by the fit. Under these circumstances, the behaviour of the contact terms cannot be reliably extrapolated to larger values of $\bar{m}_{D}$ consistent with the $T_{bc}$ case. Thus, at the present stage we cannot make a definite quantitative conclusion concerning the nature of the $T_{bc}$. Direct lattice studies of this state have not reached complete consensus yet as to whether it would be a shallow bound state [25, 26] or not [27, 28], although the more recent and extensive investigations [25, 26] suggest that a shallow bound state is more likely. Meanwhile, we notice that our qualitative inference of $T_{bc}$ as a possible shallow bound state within the effective range approximation complies well with the findings of Refs. [25, 26]. This intriguing tetraquark is being experimentally searched for by LHCb.
@@ Acknowledgements.
The authors would like to thank M. Doering, M.-L. Du, A. Filin, C. Hanhart, M. Hansen, B. Kubis, L. Leskovec, B. Mevlja, M. Mikhasenko, and A. Milstein for valuable discussions. A.N. and S.P. are supported by the Slovenian Research Agency (research core Funding No. P1-0035). A.N. also acknowledges support from the CAS President’s International Fellowship Initiative (PIFI) (Grant No. 2024PVA0004). M.P. gratefully acknowledges support from the Department of Science and Technology, India, SERB Start-up Research Grant No. SRG/2023/001235 and Department of Atomic Energy, India. We thank our colleagues in CLS for the joint effort in the generation of the gauge field ensembles which form a basis for the computation. We use the multigrid solver of Refs. [56, 57, 58, 59] for the inversion of the Dirac operator. Our code implementing distillation is written within the framework of the Chroma software package [60]. The simulations were performed on the Regensburg Athene2 cluster. We thank the authors of Ref. [61] for making the TwoHadronsInBox package public. We acknowledge the HPC RIVR consortium (www.hpc-rivr.si) and EuroHPC JU (eurohpc-ju.europa.eu) for funding this research by providing computing resources of the HPC system Vega at the Institute of Information Science (www.izum.si).
@@ Appendix A Discussion of the $DD^{*}$ interaction potential
[20]∗ 0.348 −$-3.23\pm 0.39$ $6.37\pm 1.48$−$\simeq\leavevmode\nobreak\ -15\pm 19\leavevmode\nobreak\ i$$-2.00\pm 0.47$ 0
@@ Appendix B Approach to solving Lippmann–Schwinger equation (21)
Then we find (with the subscripts I and II for the first and second Riemann sheet, respectively) that, across the unitary cut,